### v1

# What Do Latent Action Models Actually Learn?

Abstract
Latent action models (LAMs) aim to learn action-relevant changes from unlabeled videos by compressing changes between frames as latents. However, differences between video frames can be caused by controllable changes as well as exogenous noise, leading to an important concern – do latents capture the changes caused by actions or irrelevant noise? This paper studies this issue analytically, presenting a linear model that encapsulates the essence of LAM learning, while being tractable. This provides several insights, including connections between LAM and principal component analysis (PCA), desiderata of the data-generating policy, and justification of strategies to encourage learning controllable changes using data augmentation, data cleaning, and auxiliary action-prediction. We also provide illustrative results based on numerical simulation, shedding light on the specific structure of observations, actions, and noise in data that influence LAM learning.
1 Introduction
Latent action models (LAMs) aim to infer controllable action changes from streams of image observations in an unsupervised manner (Rybkin et al., 2018; Menapace et al., 2021). This is valuable because action-labeled data is typically expensive to source, while unlabeled videos are abundant. Hence, it offers a route for embodied AI systems to learn from large unlabeled datasets, for instance using the inferred latent actions as targets for pre-training a policy, while a small amount of labeled data can be used to learn a mapping from latent to real action controls (Ye et al., 2024). This has proven effective in learning from videos of 2D video games, robotics, and even broadcast tennis footage (Menapace et al., 2021; Schmidt and Jiang, 2023; Bruce et al., 2024; Chen et al., 2024b; Ye et al., 2024; Sun et al., 2024; Cui et al., 2025; Gao et al., 2025).
The success of such LAM-based recipes relies on the inferred latent action labels mapping to the real control action signals of interest. However, there is concern that this may not always be the case. For example, there is an intuition that LAM’s inferred latent ‘actions’ simply compress differences between consecutive frames, even when control actions are not the cause of those differences (McCarthy et al., 2024). As such, LAMs may only succeed in domains where the cause of changes between observations can be fully attributed to the control action. Another example is that, although bottleneck is widely used in different LAMs, whether it necessary or whether it can eliminate the dependence of the latent on the observation is unclear (Schmidt and Jiang, 2023).
Figure 1: Linear LAM is an abstraction of the LAMs used in previous work. Inputting consecutive observation pairs $(\boldsymbol{o},\boldsymbol{o}^{\prime})$, the LAMs output the second observation via a reconstruction loss, $\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}$. An information bottleneck tries to stop the direct copying of $\boldsymbol{o}^{\prime}$, with the expectation the latent $\boldsymbol{z}$ will correspond to the control action $\mathbf{a}$. Linear LAM captures the essence of LAM training whilst being analytically tractable. The diagrams of previous LAMs are copied from their original papers: LAPO (Schmidt and Jiang, 2023), LAPA (Ye et al., 2024), Moto (Chen et al., 2024c), Genie (Bruce et al., 2024), AdaWorld (Gao et al., 2025), and Go-1 (AgiBot-World, 2025).
To elucidate LAMs, this paper conducts a theoretical analysis of a linear version of LAM. By retaining the architecture of recent LAM models, but swapping deep neural network components with simpler linear layers, we are able to preserve the fundamental challenge of LAM training but in an analytically tractable form. Our analysis of linear LAM provides precise insights into, e.g., when inferred latent actions capture true control signals, and when they don’t, and what information is captured by different components within LAM. Surprisingly, our analysis also reveals additional issues not currently known to the LAM community – related to the over-parametrization property of LAM and the randomness of data-generation policy. Finally, we propose and study potential solutions to ameliorate these issues – data augmentation and predicting action as an auxiliary task.
Concretely, this paper makes several key contributions.
  1. 1.
    (Section 3) We present linear LAM, a framework that captures the essence of LAM in practice while being tractable.
  2. 2.
    (Section 4.1) We show linear LAM reduces exactly to principal component analysis (PCA) on a mixture of controllable changes and exogenous noise, under certain assumptions. Our analysis justifies the general LAM framework by showing that LAM is possible to capture the true action without accessing action labels.
  3. 3.
    (Section 4.2) Correlations between observations and actions in data reduces LAM’s focus on the controllable changes. This suggests that higher randomness in data-generating policies benefits LAM’s learning.
  4. 4.
    (Section 4.3) We validate that performing data augmentation during LAM training can mitigate the over-parametrization issue and thus improve the semantics of the latent.
  5. 5.
    (Section 4.4) We find that adding an action-prediction head encourages LAM to prioritize the learning of controllable changes for the latent.
2 Related Work
A similar topic with longer history is learning action representations in reinforcement learning (RL). For instance, PG-RA (Chandak et al., 2019) clusters actions based on the similarity of their impact on the state to improve generalization in the action space. LASER (Allshire et al., 2021) learns latent action through an encoder-decoder architecture to reconstruct real actions, resulting in higher learning efficiency in RL. TAP (Jiang et al., 2022) learns the latent action that can help to reconstruct the full trajectory (with state, action, and reward) condition on the state. EAR (Hua et al., 2022) finds that the latent task embedding resulting from the training of multi-task policies emerges to be good action representation with a gometrically and semantically meaningful structure. AD3 (Wang et al., 2024) adopts a inverse dynamics model and a forward dynamics model to extract latent action, similar to popular LAMs, but conditions these models on the real actions.
Different from the above papers that learn action representation based on real actions, our paper focuses on the model where the latent action is learned without access to the real action labels. Without the need for action labels during training, it is advantageous as it allows leveraging Internet-scale video datasets (Miech et al., 2019; Chen et al., 2024a; Pei et al., 2025; Wang et al., 2023) in the pre-training stage, which contain rich expert demonstrations but lack action labels. While high-quality robotic datasets with action annotations exist (Vuong et al., 2023; Fang et al., 2023; Khazatsky et al., 2024; AgiBot-World, 2025), they remain limited in scale. Such datasets can be integrated into the semi-supervised learning framework to extract latent actions (Nikulin et al., 2025) or the policy fine-tuning stage (Schmidt and Jiang, 2023; Ye et al., 2024). Although some researchers try to extract pre-defined actions from observations using computer vision techniques (Mendonca et al., 2023), latent action models (LAMs) can extract latent actions to represent more general transformation.
Information bottleneck or auto-encoder is adopted as the core technique in training LAMs in an unsupervised learning manner since Rybkin et al. (2018). ILPO (Edwards et al., 2019) learns latent actions along with the training of a policy that outputs the latent action and a world model that conditions on the latent action. Menapace et al. (2021) proposes a probabilistic action network that extracts a discrete action label and a continuous action variability embedding from consecutive observations. This network is trained jointly with an action decoder to generate video controlled by extracted actions. While these work adopts bottleneck in the learning of latent actions, their training losses are complicated by involving policy learning or recurrent networks. LAPO (Schmidt and Jiang, 2023) proposes that LAM design with a inverse dynamics model that extract the latent action and a forward dynamics model that reconstruct the next observation based on the latent action. Starting from this paper, many papers follow quite similar architectures to LAPO but target for a discrete latent action space such as FICC (Ye et al., 2022), Genie (Bruce et al., 2024), LAPA (Ye et al., 2024), Moto (Chen et al., 2024c), IGOR (Chen et al., 2024b), Go-1 (AgiBot-World, 2025), and GR001T N1 (Nvidia, 2025). However, there is a debate whether discrete latent actions are better than continuous latent actions. For example, Nikulin et al. (2025) finds that using continuous latent actions with larger bottleneck can not only result in better predictability on real actions but also lead to better performance for downstream policies. We consider continuous latent action in our paper and leave the analysis on vector quantized latent action in future work. In contrast to the popularity of LAM in recent foundation models for embodied AI, issues around the objective, learnability, and robustness of LAM have received limited attention. Consequently, our paper aims to elucidate these questions.
3 Setup
This section broadly introduces the problem setting, goal and model used in recent LAM work in practice. We then more formally detail these for linear model to be used in subsequent analysis. At last, we outline the details of our simulation setup to be used in support of our later analysis.
3.1 LAMs in Practice
Setting. The setting tackled by recent LAM work assumes access to a large dataset of pairs of observations and next observations $\mathcal{D}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i})\}_{i=1}^{N}$ and only a small subset of it has action labels $\mathcal{D}_{a}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i},\boldsymbol{a}_{i})\}_{i=1}^{N_{a}}$ , with $\lambda:=|\mathcal{D}_{a}|/|\mathcal{D}|\ll 1$. Note we drop the $i$ subscript when we do not need to refer to specific samples. In practice, $\boldsymbol{o}$ and $\boldsymbol{o}^{\prime}$ could be feature vectors extracted from the original observations (Cui et al., 2025), and $\boldsymbol{o}$ could be a stack of historical observations to account for partial observability (Bruce et al., 2024). The two datasets are assumed to come from the same distribution in most LAM work (Menapace et al., 2021; Bruce et al., 2024; Ye et al., 2024) [^1] Certain work explores mismatch case, e.g., $\mathcal{D}$ is human videos and $\mathcal{D}_{a}$ is robotics data (Chen et al., 2024b)..
Model. Recent LAM designs (Schmidt and Jiang, 2023; Bruce et al., 2024; Ye et al., 2024; Chen et al., 2024b; AgiBot-World, 2025) (see Figure 1) take a pair of consecutive observations $(\boldsymbol{o},\boldsymbol{o}^{\prime})$ as input, and output a latent $\boldsymbol{z}$ as well as a prediction of the next observation $\hat{\boldsymbol{o}}^{\prime}$. (We avoid calling $\boldsymbol{z}$ ‘latent action’ to avoid confusion with the real action.)
The LAM is typically decomposed into an inverse dynamics model (IDM) and forward dynamics model (FDM), implemented as deep neural networks, and trained via a reconstruction loss,
$\boldsymbol{z}=\psi_{\text{IDM}}(\boldsymbol{o},\boldsymbol{o}^{\prime}),\;\;\;\hat{\boldsymbol{o}}^{\prime}=\psi_{\text{FDM}}(\boldsymbol{o},\boldsymbol{z}),\;\;\mathcal{L}:=\min_{\psi_{\text{IDM}},\psi_{\text{FDM}}}\mathbb{E}_{\mathcal{D}}\big{[}\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert^{2}_{2}\big{]}.$ (1)
where $\psi_{\text{IDM}}$ contains a bottleneck so the dimension of $\boldsymbol{z}$ is smaller than that of $\boldsymbol{o}$.
Use cases. We identify two primary downstream use cases of LAMs.
  • The LAM latents can be used as input to a world model $\hat{T}$ trained to generate future frames, $\hat{T}(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{z})$ (Sun et al., 2024; Chen et al., 2024b; Bruce et al., 2024; Menapace et al., 2021). These world models can generate higher-quality frames than the FDM in LAM, but have been used to only qualitatively interpret the meaning of the learned latents.
  • The LAM latents can be used as labels in the pre-training of a latent policy $\pi_{\text{latent}}(\hat{\boldsymbol{z}}|\boldsymbol{o})$ (similar to behavior cloning on actions). This policy may later be mapped to real actions, $\pi_{\text{map}}(\hat{\boldsymbol{a}}|\hat{\boldsymbol{z}})$  (Bruce et al., 2024; Schmidt and Jiang, 2023), or be followed by a fine-tuning phase on real actions (Schmidt and Jiang, 2023; Ye et al., 2024; Chen et al., 2024b).
In both use cases, it is hoped that the learned latents can be aligned with the true actions as closely as possible. As in Schmidt and Jiang (2023), “our hypothesis is … [this] may allow us to learn a latent action representation with a structure closely corresponding to the true action space”.
3.2 Linear LAM
Setting. We conduct analysis in the controlled Markov process (CMP) framework (Puterman, 2014) with $(\mathcal{O},\mathcal{A},T)$. At a given timestep, the agent receives observation $\boldsymbol{o}\in\mathcal{O}$ and takes action $\boldsymbol{a}\in\mathcal{A}$. The transition function describes the probability distribution over the next observation, $T(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{a})$.
We consider vector observations, $\mathcal{O}=\mathbb{R}^{d_{o}}$ (which could be thought of as image observations processed with a pre-trained image encoder as in Cui et al. (2025); Chen et al. (2024b)). The action space $\mathcal{A}=\mathbb{R}^{d_{a}}$ has lower dimensionality than $\mathcal{O}$, i.e., $d_{a}\ll d_{o}$ [^2] Discrete actions can be represented as one-hot vectors within $\mathbb{R}^{d_{a}}$. While most of analysis (except that on bottleneck) should also apply to discrete actions, we confine our analysis in continuous actions.. These actions are mapped to create controllable changes in observation space via an action effect matrix $X\in\mathbb{R}^{d_{o}\times d_{a}}$, $\boldsymbol{q}=X\boldsymbol{a}$.
We assume linear LAM cannot access the action $\boldsymbol{a}$ during the training, but we will make use of it to evaluate the learned LAM in simulations. We choose the transition function to be an additive combination of controllable changes $\boldsymbol{q}\in\mathbb{R}^{d}$ (the state changes caused by the ego agent’s actions) and exogenous noise $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ (representing environmental stochasticity or other agent’s actions), i.e. with the next state generated via, $\boldsymbol{o}^{\prime}=\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}$.
Model. For linear LAM, the IDM and FDM consist of linear mappings. Summarized in Figure 2, the IDM and FDM are given by,
$\boldsymbol{z}$ $=\psi_{\text{IDM}}^{\text{Linear}}(\boldsymbol{o},\boldsymbol{o}^{\prime})\coloneqq C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$ (2)
$\hat{\boldsymbol{o}}^{\prime}$ $=\psi_{\text{FDM}}^{\text{Linear}}(\boldsymbol{o},\boldsymbol{z})\coloneqq A\boldsymbol{o}+B\boldsymbol{z}=A\boldsymbol{o}+B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime}),$ (3)
where $\boldsymbol{z}\in\mathcal{Z}=\mathbb{R}^{d_{z}}$ with $d_{z}\ll d_{o}$ is the latent. All matrices are learnable parameters, including FDM parameters $A\in\mathbb{R}^{d_{o}\times d_{o}}$, $B\in\mathbb{R}^{d_{o}\times d_{z}}$, and IDM parameters $C,D\in\mathbb{R}^{d_{z}\times d_{o}}$.
As for practical LAM (Eq. 1), the linear LAM is trained via a reconstruction loss,
$\mathcal{L}(A,B,C,D)=\mathbb{E}\left[\lVert\hat{\boldsymbol{o}}^{\prime}_{i}-\boldsymbol{o}^{\prime}_{i}\rVert_{2}^{2}\right].$ (4)
Figure 2: Overview of linear LAM. Grey blocks represent learnable parameter matrices, giving rise to the predictive model $\hat{\boldsymbol{o}}^{\prime}=A\boldsymbol{o}+B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})$. The green parts illustrate linear LAM with data augmentation to reduce the amount of information the latent contains about the observation (in Section 4.3). The red parts illustrate linear LAM with auxiliary action prediction head to encourage the latent to focus on the controllable actions and suppress the noise signal (in Section 4.4).
Goal. Following our discussion of use-cases of LAM cases in Section 3.1, we interpret the desired latents to contain as much information as possible about the real action, while minimizing the amount of information about the exogenous noise and the observation. This can be formalized as,
LAM $\text{LAM Objective}\coloneqq\max_{\boldsymbol{z}}\left[\mathcal{I}(\boldsymbol{z};\boldsymbol{a})-\mathcal{I}(\boldsymbol{z};\boldsymbol{\epsilon})-\mathcal{I}(\boldsymbol{z};\boldsymbol{o})\right],$ (5)
where $\mathcal{I}$ is mutual information. Hence, an ideal latent should contain all information about the action label, and no other parts of the environment, consistent with the argument from previous work – “biasing [the latent] towards simpler representations is likely preferrable” (Schmidt and Jiang, 2023).
Whilst mutual information is typically a challenging quantity to measure, in our linear model we can approximate this straightforwardly. After training the linear LAM via Eq. 4 and freezing the parameters, we fit three additional linear layers predicting $(\boldsymbol{q},\boldsymbol{\epsilon},\boldsymbol{o})$ from $\boldsymbol{z}$, to give $(\hat{\boldsymbol{q}},\hat{\boldsymbol{\epsilon}},\hat{\boldsymbol{o}})$ respectively. Note that, under the assumption that the variables $\mathbf{x}$ and $\mathbf{y}$ are independent multivariate Gaussian variables, the mutual information $\mathcal{I}(\mathbf{x},\mathbf{y})=-\dfrac{1}{2}\log\left(\frac{\lVert\hat{\mathbf{y}}-\mathbf{y}\rVert_{2}^{2}}{\mathbb{V}\text{ar}(\mathbf{y})}\right)$ where $\hat{\mathbf{y}}$ is the least squares estimate (LSE) of $\mathbf{y}$ based on $\mathbf{x}$, and $\mathbb{V}\text{ar}(\cdot)$ indicates the total variance (see e.g., Chapter 8 of (Johnson et al., 2002)) of a multivariate random variable. We define an evaluation metric that capture the objective of training linear LAM,
[…]
Note that we get rid of the $\log(\cdot)$ function wrapped around the MSE loss to better calculate the values for this objective since the value range for the mutual information $[0,+\infty)$ is unbounded. This objective is maximized when $\boldsymbol{z}$ perfectly predicts $\boldsymbol{q}$ while containing no information about $\boldsymbol{\epsilon}$ and $\boldsymbol{o}$, resulting in the optimal value for LLO equal to $0+1+1=2$.
See the discussion on the gap between linear LAM and practical LAM in Appendix A.
4 Analysis
At times, we will present numerical simulations of linear LAM to visually communicate later analysis. See the details of simulation in Appendix B and the code in supplementary.
4.1 Analysis 1: Linear LAM is PCA
This section first shows that training linear LAM is equivalent to performing PCA on the mixture of controllable changes and exogenous noise. This requires that the controllable changes and exogenous noise $\boldsymbol{q},\boldsymbol{\epsilon}$ are uncorrelated with the observation $\boldsymbol{o}$ (Section 4.2 relaxes this assumption).
We discuss the insight that this connection provides, followed by an analysis of several important cases covering specific settings of controllable changes and exogenous noise
Proposition 4.1 (Linear LAM is PCA).
Under the linear LAM model and setup defined in Section 3.2, and additionally assuming $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\mathbf{0}$, the objective of linear LAM is equivalent to performing PCA on a mixture of controllable changes $\boldsymbol{q}$ and exogenous noise $\boldsymbol{\epsilon}$,
$\mathcal{L}=\mathbb{E}\left[\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}\right]$ (7)
Note, $BD$ is a low-rank matrix to capture the main components of $\boldsymbol{q}+\boldsymbol{\epsilon}$. See proof in Appendix C.1.
Given that we have transformed the optimization problem of linear LAM into a PCA problem (which can also be viewed as a linear auto-encoder), PCA’s property applies to linear LAM.
Proposition 4.2 (Linear LAM tries to capture $\boldsymbol{q}+\boldsymbol{\epsilon}$).
Denote the covariance matrix of $\boldsymbol{q}+\boldsymbol{\epsilon}$ as $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=\mathbb{E}[(\boldsymbol{q}+\boldsymbol{\epsilon})(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]$ and its eigenvalue decomposition $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=U\Lambda U^{T}$ with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d_{o}}$. Under the same conditions of Proposition 4.1, $\mathcal{L}$ is optimized when
  1. 1.
    $B=U_{d_{z}}$ spans the subspace of the top $d_{z}$ principal components of $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}$ where $U_{d_{z}}$ contains the first $d_{z}$ columns of $U$;
  2. 2.
    $D=U_{d_{z}}^{T}$ such that the reconstruction $BD(\boldsymbol{q}+\boldsymbol{\epsilon})=U_{d_{z}}U_{d_{z}}^{T}(\boldsymbol{q}+\boldsymbol{\epsilon})$ projects $\boldsymbol{q}+\boldsymbol{\epsilon}$ onto the subspace spanned by $U_{d_{z}}$.
The minimum loss is given by the sum of the eigenvalues of corresponding to the discarded principal components $\mathcal{L}^{*}=\sum_{i=d_{z}+1}^{d_{o}}\lambda_{i}$.
This proposition is equivalent to the Eckart–Young–Mirsky theorem (Eckart and Young, 1936) whose proof can be found in Chapter 2.4 of Golub and Van Loan (2013).
Over-parameterization issue. In fact, linear LAM is over-parameterized, resulting in the solution of $(A,B,C,D)$ for loss minimization being not unique. For example, the latent $\boldsymbol{z}=C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$ is not necessarily confined to capturing information solely from $\boldsymbol{q}+\boldsymbol{\epsilon}$. Actually, it can be an arbitrary linear combination of the information from both $\boldsymbol{q}+\boldsymbol{\epsilon}$ and $\boldsymbol{o}$ since $A$ will later compensate to attenuate $\boldsymbol{o}$’s information in $\boldsymbol{z}$. To be more specific, there exists a family of solutions $B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=(\boldsymbol{q}+\boldsymbol{\epsilon})+\alpha\boldsymbol{o}$ and $A=(1-\alpha)I$ for any $\alpha$ such that $\hat{\boldsymbol{o}}^{\prime}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})+A\boldsymbol{o}=(\boldsymbol{q}+\boldsymbol{\epsilon})+(1-\alpha)\boldsymbol{o}+\alpha\boldsymbol{o}=\boldsymbol{o}^{\prime}$.
We will revisit this issue in Section 4.3 and use data augmentation to handle this over-parameterization issue. For the ease of analysis before adopting this technique, we predict $(\hat{\boldsymbol{q}},\hat{\boldsymbol{\epsilon}},\hat{\boldsymbol{o}})$ from a surrogate latent $\tilde{\boldsymbol{z}}:=B^{-1}(\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o})$ when calculating LLO to get around this issue. This surrogate latent is the same as the original one when $A=I$, which is the case when data augmentation is adopted. Hence, we have $C=-D$ which indicates that the semantic meaning of the latent $\boldsymbol{z}=C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$ remains the same (no movement) when $\boldsymbol{o}^{\prime}=\boldsymbol{o}$ across different observations.
Figure 3: LLO (Linear LAM objective) defined (6) measured when 1) $\boldsymbol{\epsilon}=0$ (left), 2) $\boldsymbol{\epsilon}$ is i.i.d. noise (middle), and 3) $\boldsymbol{\epsilon}$ contains the effect of other agents (right). Action MSE, noise MSE, and observation MSE are the three terms in (6) respectively. We set real action dimension $d_{a}$ and exogenous action dimension $d_{b}$ to $8$ in this experiment unless otherwise stated, and ensure $\boldsymbol{q}$ has unit variance.
Case 1: When $\boldsymbol{\epsilon}=0$. Without the interplay of the exogenous noise $\boldsymbol{\epsilon}$, linear LAM can capture the true action $\boldsymbol{a}$ in the latent $\boldsymbol{z}$ even if it does not have access to $\boldsymbol{a}$, when the bottleneck is set properly.
Considering that $\boldsymbol{\epsilon}=0$ and $\boldsymbol{q}=X\boldsymbol{a}$, the covariance matrix of $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=\Sigma_{\boldsymbol{q}}=\mathbb{E}[X\boldsymbol{a}\boldsymbol{a}^{T}X]$ only has $d_{a}$ non-zero eigenvalues. Following Proposition 4.2, we can make the following conclusions.
  • When $d_{z}\geq d_{a}$, the capacity of the latent is large enough to capture all the information about the controllable change $\boldsymbol{q}$, the minimum loss $\mathcal{L}^{*}=0$, and LLO achieves the optimal.
  • Specifically, when $d_{z}=d_{a}$, the subspace spanned by the columns of $B=U_{d_{a}}$ is the same to the subspace spanned by the columns of action effect matrix $X$ (by noting that $U_{d_{a}}\Lambda U_{d_{a}}^{T}=\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=X\mathbb{E}[\boldsymbol{a}\boldsymbol{a}^{T}]X^{T}$ ). In this case, the learned latent $\boldsymbol{z}$ (whose effect is interpreted by $B$) fully captures the information of $\boldsymbol{a}$ (whose effect is $X$), and LLO is maximized. In other words, for this ideal case linear LAM’s latent perfectly captures the information of the true action $\boldsymbol{a}$ without access to it.
  • When $d_{z}<d_{a}$, $\mathcal{L}^{*}>0$ and this linear auto-encoder captures the first $d_{z}$ components in $\boldsymbol{q}$.
We illustrate how LLO varies across different $d_{a}$ and $d_{z}$ through numerical simulation in Figure 3 (left). The simulation validates that linear LAM is optimized in terms of LLO when $d_{z}\geq d_{a}$.
Case 2: When $\boldsymbol{\epsilon}$ is i.i.d. noise. We start analyze the noise from the i.i.d. (independent and identically distributed) assumption, which is an approximation for many real noises such as the noise from the sensors or the image encoders. Now we assume $\boldsymbol{\epsilon}$ is i.i.d. noise with zero-mean $\mathbb{E}[\boldsymbol{\epsilon}]=0$ and covariance $\mathbb{E}[\boldsymbol{\epsilon}\boldsymbol{\epsilon}^{T}]=\sigma_{iid}^{2}I$. Considering that $\boldsymbol{q}$ and $\boldsymbol{\epsilon}$ are independent, the covariance matrix $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}$ can be eigenvalue decomposed as $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=U\Gamma_{0}U^{T}=U_{0}(\Lambda_{0}+\sigma_{iid}^{2}I)U_{0}^{T}$ where $U_{0}$ and $\Lambda_{0}$ are the eigenvectors and eigenvalues of $\Sigma_{\boldsymbol{q}}$ respectively (i.e., when $\boldsymbol{\epsilon}=0$). Combining the results in Proposition 4.2, we conclude that, when there is i.i.d. noise, 1) the FDM parameter $B$ (and therefore the semantics of the latent since $B$ interpret the latent) remains the same, and 2) the loss increases since the eigenvalues increase. This conclusion is consistent with the conclusion on the robustness of PCA (Anderson, 1963; Johnstone, 2001).
We present a simulation results on how LLO changes across different levels of the i.i.d. Gausion noise in Figure 3 (middle). We observe that linear LAM is robust to i.i.d. noise - even when the noise level $\sigma_{iid}=0.5$ (when the noise intensity is half that of the signal), linear LAM still succeeds with negligible gap to the optimal case.
Case 3: When $\boldsymbol{\epsilon}$ contains the effect of other agents. In many practical datasets (such as Ego4d (Grauman et al., 2022)), the change between two observations is the effect of not only the controllable action but also other agents. For example, there may be another person walking in the background or camera shaking.
Here, we assume $\boldsymbol{\epsilon}=Y\mathbf{b}$ where $Y\in\mathbb{R}^{d_{o}\times d_{b}}$ is the exogenous effect matrix and $\mathbf{b}$ represents the action taken by other agents. In this case, we can conduct an analysis similar to that of Case 1, and conclude that linear LAM will learn to capture the effects with top variances, no matter they result from the controllable action or other agents. Moreover, when the columns in $Y$ are not orthogonal to that of $X$, the FDM parameter $B$ will be impacted by the exogenous noise.
We show the simulation results in Figure 3 (right) to illustrate which part of information (the controllable $\boldsymbol{q}$ or the noise $\boldsymbol{\epsilon}$) enters the latent when the latent dimension $d_{z}$ increases gradually. We can observe that: 1) When the variance of exogenous noise is smaller than $\boldsymbol{q}$ (i.e., $\sigma_{exo}=0.5$), linear LAM learns to fit the controllable part first, and $d_{z}=d_{a}$ is still the optimal configuration. 2) When the variance of noise exceeds that the signal $\boldsymbol{q}$ (i.e., $\sigma_{exo}=2.0$), linear LAM will first fit the noise, and the latent fails to exclude the information of the noise no matter how we set $d_{z}$. In this case, one possible solution is to clean training data to remove significant noises (such as shaking clips).
4.2 Analysis 2: Effects of Data Collection Policy
Till now, we consider the cases where both the controllable changes and the exogenous noise are independent of the observation, i.e., $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\mathbf{0}$. Unfortunately, this is usually not true for practical cases where LAM is trained using data collected by expert policies, e.g., many robot datasets are collected via teleoperation by humans. Within such datasets, the observation is correlated with the action $\boldsymbol{a}$, and thus the controllable change $\boldsymbol{q}$, resulting in $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]\neq\mathbf{0}$. Therefore, we dive into this scenario to see what is the effect of the data collection policy on linear LAM in this part.
Specifically, we assume the data is collected by a policy that generate action via $\boldsymbol{a}=\Pi_{d}\boldsymbol{o}+\boldsymbol{\pi}_{s}$ where $\Pi_{d}\in\mathbb{R}^{d_{a}\times d_{o}}$ represents the deterministic part of the policy and $\mathbf{\pi}_{s}\in\mathbb{R}^{d_{a}}$ represents the stochastic part of the policy. We assume that $\boldsymbol{o}$, $\boldsymbol{\epsilon}$, and $\boldsymbol{\pi}_{s}$ are uncorrelated.
In this case, since $\mathbb{R}[\boldsymbol{o}\boldsymbol{q}^{T}]\neq 0$, the cross term in the expansion of loss (7) remains. Letting $\Sigma_{\boldsymbol{o}}:=\mathbb{E}[\boldsymbol{o}\boldsymbol{o}^{T}]$ to be the covariance matrix of the observation, we can further solve for $A$ by setting the partial derivative of the loss function w.r.t. $A$ to zero, obtaining
$A=I-(BC+BD)-\Sigma_{\boldsymbol{o}}\Pi_{d}^{T}X^{T}(BD-I)^{T}\Sigma_{\boldsymbol{o}}^{-1}.$ (8)
This differs from the previous result for the random policy where $A=I-(BC+BD)$ by the last term $\Sigma_{\boldsymbol{o}}\Pi_{d}^{T}X^{T}(BD-I)^{T}\Sigma_{\boldsymbol{o}}^{-1}$. The intuitive interpretation for this term is that $A$ can capture the changes caused by the deterministic part of the policy if these changes are not captured by the latent (noting that $BD-I$ can project the vectors onto the orthogonal complement of the column space of $B$). In other words, the latent may not capture the effect of each real action. Instead, the FDM would capture the effect of the deterministic part of the policy, whereas the latent would capture the remaining stochastic part.
In our simulation, we control the randomness of the policy using a coefficient $\chi$, and let $\boldsymbol{a}=\chi\Pi_{d}\boldsymbol{o}+(1-\chi)\boldsymbol{\pi}_{s}$ where $\Pi_{d}$ is a random projection matrix and $\pi_{s}$ is a standard Gaussian vector. The simulation results in Figure 4 (left) indicate that the more deterministic the data collection policy is, the less information about $\boldsymbol{a}$ is captured by $\boldsymbol{z}$. Therefore, the analysis in this section suggest us to use the data collected by the policies with higher randomness whenever possible. This may also indicate that, for LAM training, the trajectories collected by expert policies may not be enough, and more exploratory trajectories are needed.
Figure 4: LLO (Linear LAM objective) defined (6) measured when 1) data is collected by policies of different deterministic levels (left), 2) linear LAM is trained w/ and w/o data augmentation (middle), and 3) linear LAM is trained w/ and w/o action prediction (right). We set $d_{a}=8$ in the experiments.
4.3 Analysis 3: Improvements via Data Augmentation
This section analyzes the augmentation scheme used in Chen et al. (2024b), which will turn out to resolve a key issue encountered in linear LAM (see the discussion on over-parameterization in Section 4.1), where information about the observation $\boldsymbol{o}$ enters into the latent $\boldsymbol{z}$.
The IGOR model (Chen et al., 2024b) applies one shared random crop to both inputs of the IDM, and a second shared random crop applied to the FDM and reconstruction target. The intuition is that “by using different croppings [for IDM and FDM], the model is encouraged to learn a more semantically invariant latent action”. Our subsequent analysis shows this intuition is provably correct in linear LAM, resulting in new terms to the loss that encourage removal of the latent of any semantic information about the observation. Sun et al. (2024) also apply a similar scheme, with a single action preserving crop applied to the IDM which makes it harder “to copy the appearance information directly from the [IDM]”.
Data augmentation in linear LAM. We extend our linear LAM setup to include a data augmentation operator $\operatorname{Aug}_{i}[\boldsymbol{o}]\coloneqq\boldsymbol{o}+\boldsymbol{\kappa}_{i}$ that takes an observation as input and adding some random vector $\boldsymbol{\kappa}\in\mathbb{R}^{d}_{o}$. $\operatorname{Aug}_{i}[\cdot]$ applies the same $i$-th operator to different variables, say $\operatorname{Aug}_{1}[\boldsymbol{o}]$ and $\operatorname{Aug}_{1}[\boldsymbol{o}^{\prime}]$ will apply a consistent random variable $\boldsymbol{\kappa}_{1}$ to each observation.
Proposition 4.3 (Data augmentation addresses over-parameterization).
With data augmentation,
$\boldsymbol{z}$ $=\psi_{\text{IDM}}^{\text{Linear}}(\operatorname{Aug}_{1}[\boldsymbol{o}],\operatorname{Aug}_{1}[\boldsymbol{o}^{\prime}]_{1})\coloneqq C(\boldsymbol{o}+\boldsymbol{\kappa}_{1})+D(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa}_{1})$ (9)
$\hat{\boldsymbol{o}}^{\prime}$ $=\psi_{\text{FDM}}^{\text{Linear}}(\operatorname{Aug}_{2}[\boldsymbol{o}],\boldsymbol{z})\coloneqq A(\boldsymbol{o}+\boldsymbol{\kappa}_{2})+B\boldsymbol{z}.$ (10)
and assuming $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\mathbf{0}$, optimizing the loss defined in (1) results in $A=I$ and $C+D=0$.
We provide the proof in Appendix C.2. What does it mean for these terms to encourage $A=I$ and $C+D=0$? If, $A$ is the identity, all observation information flows directly through this matrix, and $\boldsymbol{z}$ need not carry any additional information about $\boldsymbol{o}$, under the additive assumption on the dynamics. With similar effect, setting $C+D=0$ immediately cancels out $\boldsymbol{o}$ information in $\boldsymbol{z}$.
$\boldsymbol{z}=C(\boldsymbol{o}+\boldsymbol{\kappa}_{1})+D(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa}_{1})=-D(\boldsymbol{o}+\boldsymbol{\kappa}_{1})+D(\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}+\boldsymbol{\kappa}_{1})=D(\boldsymbol{q}+\boldsymbol{\epsilon})$ (11)
This condition also hints semantic consistency – the semantic meaning of the latent the frame does not change from $\boldsymbol{o}$ to $\boldsymbol{o}^{\prime}$ will be consistent across different observations. Hence, this augmentation scheme is a mechanism to remove information about $\boldsymbol{o}$ from $\boldsymbol{z}$, explicitly minimizing $\mathcal{I}(\boldsymbol{z};\boldsymbol{o})$ in the original objective (5). In this way, data augmentation results in latents with better semantics.
Simulation. We show the effect of data augmentation under different noise level in Figure 4 (middle) The results indicate the by data augmentation can improve the semantics of the latent $\boldsymbol{z}$ by addressing the over-parameterization issue. Note that, starting here, the simulation calculate LLO based on the true latent $\boldsymbol{z}$ instead of the pseudo-latent. In this way, even when there is no noise $\sigma_{iid}=0$, the learned latent $\boldsymbol{z}$ cannot perfectly predict the real action $\boldsymbol{a}$ without involving other information. However, adding in the augmentation vector with 10% variance (compared with that of observation $\boldsymbol{o}$), can greatly improve the learning of linear LAM.
Designing data augmentation. In our linear setting, a natural design for data augmentation is to have $\boldsymbol{\kappa}$ i.i.d. across its elements. This design is based on our knowledge that the semantic meaning for the frame change should be invariant to this data augmentation, i.e., $\boldsymbol{o}^{\prime}-\boldsymbol{o}=(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa})-(\boldsymbol{o}+\boldsymbol{\kappa})$, since we know the dynamics is additive. For real images, Chen et al. (Chen et al., 2024b) adopts random crops, reflecting the prior that the semantic meaning about frame change is invariant to this way of augmentation. Further, the variance of different augmentations determines how important this term is – a larger augmentation variance enforces this constraint more strictly. This indicates that we may design a weighted mix of data augmentation that best capture the invariance property to improve the semantics of the learned latent.
4.4 Analysis 4: Improvements via Auxiliary Action Prediction
This section analyzes the setting when a small dataset of action labeled data $\mathcal{D}_{a}$ is available during training of the LAM. This can be used to help guide the latents to represent controllable changes rather than exogenous noise. Specifically we consider the a simple auxiliary loss, with latents as input, predicting the action labels when available. Nikulin et al. (2025) also provide empirically evidence to show that this is a promising strategy to avoiding focus on ‘distractors’ present in the observations.
Action prediction in linear LAM. Consider a linear prediction head $E\in\mathbb{R}^{d_{a}\times d_{z}}$ at the latent bottleneck, $\hat{\boldsymbol{a}}=E\boldsymbol{z}$, and a corresponding action reconstruction loss $\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}$, giving the overall objective, $\mathcal{L}_{a}:=\mathbb{E}_{\mathcal{D}}\left[\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}\right]+\mathbb{E}_{\mathcal{D}_{a}}\left[\lVert\hat{\boldsymbol{a}}-\boldsymbol{a}\rVert_{2}^{2}\right]$ with $\lambda:=|\mathcal{D}_{a}|/|\mathcal{D}|$.
Proposition 4.4 (Action prediction can denoise).
Following the conditions in Proposition 4.3 and assuming $\boldsymbol{\epsilon}^{T}X=0$ and $d_{z}\geq d_{a}$, optimizing $\mathcal{L}_{a}$ biases the encoder parameter perpendicular to the noise. For an artificial case where $\lambda\to+\infty$, we obtain perfect LAM with $D\boldsymbol{\epsilon}=\boldsymbol{0}$ and $B\boldsymbol{z}=\boldsymbol{q}$.
We provide the proof in Appendix C.3. Note that we assume that $\boldsymbol{\epsilon}^{T}X=0$, which means that the signal $\boldsymbol{q}=X\boldsymbol{a}$ and the noise $\boldsymbol{\epsilon}$ are disentangled. This can correspond to the case, e.g., where $\boldsymbol{\epsilon}$ is background disturbance and $\boldsymbol{q}$ is the table-top manipulation (cf. Case 3 in Section 4.1). The conclusion $D\perp\boldsymbol{\epsilon}$ indicates that noise will not enter the latent (noting that $C=-D$ in this case).
Although Proposition 4.3 considers an artificial case with learn $\lambda$, it indicates that auxiliary action prediction can bias linear LAM to denoise. In our simulation, we induce different levels of noise caused by other agents with the same setting as in Figure 3 (right), and present the result in Figure 4 (right). In contrast to previous results where $\sigma_{exo}$ fails the learning of linear LAM, using only 1% of action label in training leads to successful learning. The simulation results indicate that only a small proportion of action label in training can significantly encourage the latent to encode the real actions and denoise.
5 Conclusion
Latent action models (LAMs) have become popular in the pre-training phase of many embodied AI models. In contrast to the popularity, detailed analysis into the learnability of LAMs is missing. This paper tries to bridge the gap by proposing a linear LAM model that captures the essence of LAM but remains mathematical tractability. By analyzing linear LAM, we have made several fruitful observations, including justifying the design of LAM training loss by showing that it corresponds to PCA and can succeed in recovering the real action, highlighting in the data collection policy’s impact on the learning of LAM, and showing the functionality of data augmentation and auxiliary action prediction. We acknowledge that we should also notice the gap between linear LAM and practical LAM in terms of both the capacity of the FDM and IDM models and the linear dynamics assumed in linear LAM. Nevertheless, we believe linear LAM can be further used as a quick testbed for researchers to develop other tricks for LAM learning.
Appendix A From practical LAM to linear LAM.
Our goal in designing linear LAM was to preserve the key features of LAMs in practice, while making the resulting model as simple as possible. Here we remark on similarities and differences.
  • Function approximation. Linear LAM uses linear layers, while practical LAM uses deep neural networks. However, linear LAM is compatible with input vectors processed by other non-linear pre-trained image encoders.
  • Bottleneck. Both models have an information bottleneck between in the IDM but this is implemented in different ways. Practical LAMs usually adopt vector quantization Van Den Oord et al. (2017), while linear LAM uses a low continuous dimension $d_{z}\ll d_{o}$.
  • Additive changes. We formulate changes between observations as additions of controllable changes and exogenous noise. This additive structure provides the simplest combination of two elements which we believe are a primary concern of whether LAM’s latents actually represent actions.
  • Noise. While our model constrains the noise as additive, it does not make further assumptions. We later analyze realistic cases corresponding to real world scenarios, such as when action and noise are correlated, and action and observations are correlated.
Appendix B Simulation Set-ups
We adopt the following setting unless otherwise stated. We provide the source code in the appendix.
  • Observations. Observations are sampled from the standard normal distribution, $\boldsymbol{o}\sim\mathcal{N}(\mathbf{0},I)$ with $I$ as the identity matrix. Note that $\mathbb{V}\text{ar}(\boldsymbol{o})=1$ (i.e., each element in $\boldsymbol{o}$ has unit variance). We set the dimension of the observations $d_{o}=128$.
  • Actions. We use continuous actions also sampled from a standard normal distribution with dimensionality $d_{a}$, so $\boldsymbol{a}\sim\mathcal{N}(\mathbf{0},I)$. Note that $\mathbb{V}\text{ar}(\boldsymbol{a})=1$. We set $d_{a}=8$ unless otherwise stated.
  • Controllable changes. The action effect matrix $X$ that maps the action $\boldsymbol{a}$ to the controllable changes $\boldsymbol{q}$, is chosen as a random orthogonal matrix (using QR decomposition), subsequently normalized to ensure $\mathbb{V}\text{ar}(\boldsymbol{q})=1$.
  • i.i.d. noise. In our simulation we use isotropic Gaussian $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\epsilon}^{2}I)$ with variance $\sigma_{iid}^{2}$ as the i.i.d. noise. For this noise, $\mathbb{V}\text{ar}(\boldsymbol{\epsilon})=\sigma_{iid}^{2}$.
  • Exogenous noise. We also consider the noise induced by other agents $\boldsymbol{\epsilon}=Y\boldsymbol{b}=\sigma_{exo}Y_{0}\boldsymbol{b}$ where $\boldsymbol{b}$ is other agents’ action, $Y$ is the action effect matrix of other agents, and $Y_{0}$ is the normalized matrix of $Y_{0}$. Similarly, $Y_{0}$ is chosen as a random orthogonal matrix using QR decomposition to ensure that $\mathbb{V}\text{ar}(\boldsymbol{\epsilon})=\sigma_{exo}^{2}$.
  • Optimization. We implement our system in PyTorch, optimizing trainable parameters via stochastic gradient descent with the Adam optimizer with batch size 128. We use the default learning rate and run for 4,000 steps to ensure convergence.
  • Evaluation. Our use the quantity defined in (6) as the default evaluation metric. For the experiments that do not involve noise, we set normalized MSE for noise to $1$.
  • Data augmentation. Data augmentation is a trick that may improve the learnability of LAM mentioned in several previous papers (Chen et al., 2024b; Sun et al., 2024). We implement data augmentation by adding a Gaussian noise $\boldsymbol{\kappa}\sim\mathcal{N}(\mathbf{{0},|\boldsymbol{\kappa}|^{2}I})$ for linear LAM. Note that $\mathbb{V}\text{ar}(\boldsymbol{\kappa})=|\boldsymbol{\kappa}|^{2}$. By default data augmentation is turned off, except for Figure 4 (middle and right).
  • Action prediction. Action prediction is another trick proposed in the previous paper (Nikulin et al., 2025). We implement action prediction by predicting the true action label based on the latent with a learnable linear transformation for a small proportion of the data samples. We denote the ratio of the samples that we access their action labels as $\lambda$. We find that setting $\lambda=1\%$ is enough. By default action prediction is turned off, except for Figure 4 (right).
Appendix C Proofs
C.1 Proof of Proposition 4.1
Proof.
Expand the loss function in (4) with the definitions of $\hat{\boldsymbol{o}}^{\prime}$ (3) and $\boldsymbol{o}^{\prime}$ ($\boldsymbol{o}^{\prime}=\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}$), then rearrange (ignoring the expectation outside the RHS).
$\mathcal{L}=$ $\lVert\boldsymbol{o}^{\prime}-\hat{\boldsymbol{o}}^{\prime}\rVert_{2}^{2}$ (12)
$=$ $\lVert\boldsymbol{o}^{\prime}-(A\boldsymbol{o}+BC\boldsymbol{o}+BD\boldsymbol{o}^{\prime})\rVert_{2}^{2}$ (13)
$=$ $\lVert(\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon})-\left(A\boldsymbol{o}+BC\boldsymbol{o}+BD(\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon})\right)\rVert_{2}^{2}$ (14)
$=$ $\lVert(A+BC+BD-I)\boldsymbol{o}+(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$ (15)
$=$ $\lVert(A+BC+BD-I)\boldsymbol{o}\rVert_{2}^{2}+\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$
$\qquad\qquad\qquad\qquad\qquad\quad+2\boldsymbol{o}^{T}(A+BC+BD-I)^{T}(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})$ (16)
$=$ $\lVert(A+BC+BD-I)\boldsymbol{o}\rVert_{2}^{2}+\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$
$\qquad\qquad\qquad\qquad\qquad\quad+2\operatorname{Tr}\left((A+BC+BD-I)\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}(BD-I)\right)$ (17)
Recall that this loss $\mathcal{L}$ is found within an expectation in (4). By assumption $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\textbf{0}$ and the final term can be ignored (since both expectation and trace are additive).
$\mathcal{L}=\mathbb{E}\left[\lVert(A+BC+BD-I)\boldsymbol{o}\rVert_{2}^{2}\right]+\mathbb{E}\left[\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}\right]$ (18)
Regarding the first term, note that $A$ is full rank $d_{o}$ while $BC$ and $BD$ are of rank $d_{z}$. Since $A$ only appears in this term and it is of greater or equal rank than $BD+BD$, it’s optimal value is $A=I-B(C+D)$, setting the first term is zero. Hence we are left with the middle term.
$\mathcal{L}=\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$ (19)
Note that, under the condition of Proposition 4.1, $(A,B,C,D)$ are over-parameterized. When we adopting data augmentation in Proposition 4.3, we obtain $A=I$ and $C+D=0$, which refines how the condition $A=I-B(C+D)$ is satisfied.
C.2 Proof of Proposition 4.3
We start by expanding the loss defined in (1) with the data augmentation scheme, and unpack (ignoring the expectation outside the RHS).
$\mathcal{L}$ $=\lVert\operatorname{Aug}_{2}[\boldsymbol{o}^{\prime}]-\hat{\boldsymbol{o}}^{\prime}\rVert_{2}^{2}$ (20)
$=\lVert\boldsymbol{o}^{\prime}+\boldsymbol{\kappa}_{2}-(A(\boldsymbol{o}+\boldsymbol{\kappa}_{2})+B(C(\boldsymbol{o}+\boldsymbol{\kappa}_{1})+D(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa}_{1})))\rVert_{2}^{2}$ (21)
$=\lVert(\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon})+\boldsymbol{\kappa}_{2}-(A(\boldsymbol{o}+\boldsymbol{\kappa}_{2})+B(C(\boldsymbol{o}+\boldsymbol{\kappa}_{1})+D(\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}+\boldsymbol{\kappa}_{1})))\rVert_{2}^{2}$ (22)
$=\lVert(A+BC+BD-I)\boldsymbol{o}+(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})+(BC+BD)\boldsymbol{\kappa}_{1}+(I-A)\boldsymbol{\kappa}_{2}\rVert_{2}^{2}$ (23)
$=\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}+\lVert(A+BC+BD-I)\boldsymbol{o}\rVert_{2}^{2}$
$\qquad\qquad\qquad\qquad+\lVert B(C+D)\boldsymbol{\kappa}_{1}\rVert_{2}^{2}+\lVert(I-A)\boldsymbol{\kappa}_{2}\rVert_{2}^{2}$ (24)
Where the last line follows since $\boldsymbol{\kappa}_{1}$ and $\boldsymbol{\kappa}_{2}$ are sampled independently of all other terms and $\mathbb{E}[\boldsymbol{o}^{T}(\boldsymbol{q}+\boldsymbol{\epsilon})]=0$. Hence, we are left with the vanilla linear LAM loss in (15) plus two additional terms.
Minimizing this function can be achieved by exactly setting $A=I$ and $C=-D$, which zeros the last three terms simultaneously. In this case, the optimization problem again reduces to PCA $\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$.
Discussion. In the correlated case $\mathbb{E}[\boldsymbol{o}^{T}(\boldsymbol{q}+\boldsymbol{\epsilon})]\neq 0$, when the third term (the cross term) in (17) cannot be ignored, nevertheless there is encouragement to reach $A=I$ and $C=-D$.
C.3 Proof of Proposition 4.4
Based on the conclusion in Proposition 4.3, optimizing for the first term in $\mathcal{L}_{a}$ results in $A=I$ and $C=-D$. Since $\boldsymbol{\epsilon}^{T}X=0$, we have $\boldsymbol{q}^{T}\boldsymbol{\epsilon}=0$. We consider the case where $\mathcal{D}$ and $\mathcal{D}_{a}$ come from the same distribution and the availability of action labels are uniformly random. Therefore, we can ignore the expectation outside the RHS (for simplicity) and re-write the loss as,
$\mathcal{L}_{a}$ $=\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}+\lambda\lVert ED\left(\boldsymbol{q}+\boldsymbol{\epsilon}\right)-\boldsymbol{a}\rVert_{2}^{2}$ (25)
$=\lVert(BD-I)\boldsymbol{q}\rVert_{2}^{2}+\lVert(BD-I)\boldsymbol{\epsilon}\rVert_{2}^{2}+\lambda\lVert ED\boldsymbol{q}-\boldsymbol{a}\rVert_{2}^{2}+\lambda\lVert ED\boldsymbol{\epsilon}\rVert_{2}^{2}$ (26)
$=\lVert(BDX-X)\boldsymbol{a}\rVert_{2}^{2}+\lVert(BD-I)\boldsymbol{\epsilon}\rVert_{2}^{2}+\lambda\lVert(EDX-I_{d_{a}})\boldsymbol{a}\rVert_{2}^{2}+\lambda\lVert ED\boldsymbol{\epsilon}\rVert_{2}^{2}$ (27)
We decompose $X$ as $X=U\Sigma V^{T}$ with $U\in\mathbb{R}^{d_{o}\times d_{a}}$ (which spans the column space of $X$), $\Sigma\in\mathbb{R}^{d_{a}\times d_{a}}$, and $V\in\mathbb{R}^{d_{a}\times d_{a}}$.
We will show that, when $\lambda\to+\infty$, $D=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]$, $B=[U,*]$, and $E=[V\Sigma^{-1},*]$ minimizes $\mathcal{L}_{a}$ where * indicates arbitrary entries and * vanishes when $d_{z}=d_{a}$.
First, this solution zeros the last two terms. For the last term, $D\boldsymbol{\epsilon}=\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]\boldsymbol{\epsilon}=0$ since $\boldsymbol{\epsilon}^{T}X=0$ and $U$ spans the column space of $X$. For the third term, $EDX=[V\Sigma^{-1},*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]U\Sigma V^{T}=I_{d_{a}}$. Then, since we have determined $D$, the second term becomes $\lVert(BD-I)\boldsymbol{\epsilon}\rVert_{2}^{2}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can choose $B=[U,*]$ to zero the first term since $BDX=[U,*]\left[\begin{matrix}U^{T}\\ 0\end{matrix}\right]X=X$. In this way, we obtain the minimal loss $\mathcal{L}^{*}_{a}=\lVert\boldsymbol{\epsilon}\rVert_{2}^{2}$. We can also verify that $B\boldsymbol{z}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=BD\boldsymbol{q}=UU^{T}\boldsymbol{q}=\boldsymbol{q}$.
Though the $\lambda\to+\infty$ setting is artificial, the analysis show that a positive $\lambda$ will bias the encoder $D$ to capture less about the noise $\boldsymbol{\epsilon}$ and more about the signal $\boldsymbol{q}$ (or $\boldsymbol{a}$).